# BootsTAP: Bootstrapped Training for Tracking-Any-Point

### v1

@@ Abstract
To endow models with greater understanding of physics and motion, it is useful to enable them to perceive how solid surfaces move and deform in real scenes. This can be formalized as Tracking-Any-Point (TAP), which requires the algorithm to be able to track any point corresponding to a solid surface in a video, potentially densely in space and time. Large-scale ground-truth training data for TAP is only available in simulation, which currently has limited variety of objects and motion. In this work, we demonstrate how large-scale, unlabeled, uncurated real-world data can improve a TAP model with minimal architectural changes, using a self-supervised student-teacher setup. We demonstrate state-of-the-art performance on the TAP-Vid benchmark surpassing previous results by a wide margin: for example, TAP-Vid-DAVIS performance improves from 61.3% to 66.4%, and TAP-Vid-Kinetics from 57.2% to 61.5%.
@@ 1 Introduction
Despite impressive achievements in the vision and language capability of generalist AI systems, recent research continues to cite lack of physical and spatial reasoning as a weakness of state-of-the-art vision models [40, 49]. This limits their application in many domains like robotics, video generation, and 3D asset creation – all of which require an understanding of the complex motions and physical interactions in a scene. Tracking-Any-Point (TAP) [9] is a promising approach to represent precise motions in videos, and recent work has demonstrated compelling usage of TAP in robotics [52, 58], video generation [10], and video editing [60]. In TAP, algorithms are fed a video and a set of query points—potentially densely across the video—and must output the tracked location of these query points in the video’s other frames. If the point is not visible in a frame, the point is marked as occluded in that frame. This approach has many advantages. It is a highly general task, as correspondences for surface points are typically well-defined for solid surfaces, and it also serves as a rich source of information about the deformation and motion of objects across long time periods.
@@
The main challenge for building TAP models, however, is the lack of training data: in the real world, we must rely on manual labeling, which is arduous and imprecise [9], or on 3D sensing [1], which is only available in limited scenarios. Thus, state-of-the-art methods have relied on synthetic data [16, 63]. In this work, however, we demonstrate that unlabeled real-world videos can be used to improve point tracking, using self-consistency as a supervisory signal. In particular, we know that if the tracks are correct for a given video, then 1) spatial transformations of the video should result in an equivalent spatial transformation of the trajectories, 2) that different query points along the same trajectory should produce the same track, and 3) that non-spatial data augmentation (e.g. image compression) should not affect results. Deviations from this can be treated as an error signal for learning.
Our architecture is outlined in Figure 1. We begin with a strong “teacher” model pre-trained using supervised learning on synthetic data (in our case, a TAPIR [10] model) which serves as initialization for both a “teacher” and a “student” model. Given an unlabeled input video, we make a prediction using the teacher model, which serves as pseudo-ground-truth for the student. We then generate a second “view” of the video by applying affine transformations that vary smoothly in time, re-sampling frames to a lower resolution (while padding back to the original size), and adding JPEG corruption. We input the second view to the “student” network and use a query point sampled from the teacher’s prediction (transformed consistently with the transformation applied to the video). The student’s prediction is then transformed back into the original coordinate space. We then compute a “self-supervised loss,” which is simply TAPIR’s original loss function applied to the student predictions, using teacher’s predictions as pseudo-ground-truth. The teacher’s weights are updated by using an exponential moving average (EMA) of the student’s weights. We take steps to ensure that the teacher’s predictions used for training are more likely to be accurate than the student’s: (i) the corruptions that degrade and downsample the video are only applied to the student’s inputs, (ii) we use an EMA of the student’s weights as the teacher’s weights, which we empirically find makes the teacher’s outputs more accurate and stable, finally, (iii) we only train on points that are closer (in time) to the teacher’s query point than they are to the student’s. We show that this formulation, when applied to real-world videos (15M clips) provides a substantial boost over prior state-of-the-art across the entire TAP-Vid benchmark.
@@
  3. 3.
    We release a checkpoint on GitHub, and include model implementations in both JAX and PyTorch for the community to use, available at https://github.com/google-deepmind/tapnet.
@@ 2 Related Work
The ability to track densely-sampled points over long video sequences is a generic visual capability [44, 45]. Because this visual task provides a rich output that is well-defined independent of semantic or linguisitic categories (unlike classification, detection, and semantic segmentation), it is more generically useful and can support other visual capabilities like video editing [60], 3D estimation [54], object segmentation [39, 42], and even robotics [52, 58]. Point tracking has recently experienced a flurry of recent works including new datasets [9, 63, 1] and algorithms [18, 25, 10, 54, 37]. Current state-of-the-art works mainly train in a supervised manner, relying heavily on synthetic data [16] which has a large domain gap with the real world.
@@
A final self-supervised approach is to create synthetic pairs from real images where correspondences are known [19, 46]. Such approaches have a long history in optical flow [31, 32, 20, 38], although with mixed results, typically requiring complex training setups such as GANs [27] or connecting the student to the teacher [33] to prevent trivial solutions. They have only been applied to longer-term point tracking more recently [54, 48]. OmniMotion computes initial point tracks using RAFT [50] or TAP-Net [9] and infers a full pseudo-3D interpretation of the scene in the form of a neural network. Although this method improves point tracks compared to their initialization, it never retrains a general TAP model on the self-labeled data. Concurrent work [48], on the other hand, saves a dataset of point tracks and retrains the underlying model on them, using data augmentations similar to ours. Unlike our work, however, it trains only on relatively small, downstream datasets, and it does not update the underlying point tracks as the model improves. They report that performance quickly saturates and begins to degrade with more training iterations, which is a key limitation for scaling the method that our algorithm overcomes. We discuss the differences in more detail in the following section, after presenting our approach.
@@ 3 Method
When developing a self-training method for TAP, it is important to note that, like in optical flow, TAP has a precise, correct answer for the vast majority of query points. This is different from typical visual self-supervised learning, where the representation can be arbitrary, as long as semantically similar images have similar representations. Supervised learning on synthetic data provides a strong initial guess in many situations, but care must be taken to ensure that the self-supervised algorithm does not find “trivial shortcuts” [8] that become self-reinforcing and harm the initialization.
@@
Figure 2: Bootstrapped training for Tracking-Any-Point. The teacher TAPIR produces a pseudo-label trajectory from query point $q_{1}$ at time $t_{1}$. Video frames undergo affine transformations that vary smoothly in time and are augmented with JPEG artifacts, then fed to the student TAPIR, which predicts a trajectory from query point $\Phi(q_{2})$ at time $t_{2}$ (sampled from the teacher’s prediction, then transformed to the student video space using $\Phi$). The student trajectory is transformed back, and loss is computed against the teacher’s trajectory. Note that only trajectories where the student’s prediction on at time $t_{1}$ are close to the teacher contribute to the loss (orange circle), and only points temporally closer to the teacher’s query frame than to the student’s query frame (green/red points on the left) contribute to the loss.
@@
Our formulation relies on two facts about point tracks that are true for points on any solid, opaque surface. First, spatial transformations (e.g. affine transformations) which are applied to the video will result in equivalent spatial transformations of the point tracks (i.e. the tracks are “equivariant” under spatial transformation), while the tracks are invariant to many other factors of variation that do not move the image content (e.g. color changes, noise). Second, the algorithm should output the same track regardless of which point along the track is used as a query; mathematically, this means that each trajectory forms an equivalence class. One could imagine enforcing the desired equivariance and invariance properties using a simple Siamese-network formulation [17], where a single network is trained to output consistent predictions on two different ‘views’ of the data. However, we find that minimizing the difference between the two outputs—and backpropping both—results in predictions degrading toward trivial solutions (e.g. over-smoothing of tracks, or tracking the image boundary instead of the image contents). In fact, the model can learn to distinguish between synthetic and real data resulting in trivial solutions on the real, unlabeled data only. To prevent this, we adopt a student-teacher framework, where the student’s view of the data is made more challenging by augmentations, and the teacher does not receive gradients that may corrupt its predictions. Figure 2 shows the overall pipeline.
@@
We start with a baseline TAPIR network pre-trained on Kubric in the standard way [10]. Let $\hat{y}=\{\hat{p},\hat{o},\hat{u}\}$ be the predictions: $\hat{p}\in\mathbb{R}^{T\times 2}$ is position, $\hat{o}\in\mathbb{R}^{T}$ is an occlusion logit, and $\hat{u}\in\mathbb{R}^{T}$ is an uncertainty logit, where $T$ is the number of frames. Calling $p[t]$ and $o[t]$ the ground truth targets for frame $t$, recall that the standard TAPIR loss for a single trajectory is defined as:
@@
After pre-training, we add a few extra convolutional residual layers to the backbone, initialized to the identity [15], to absorb the extra training data. Let $\hat{y}_{\mathcal{S}}=\{\hat{p}_{\mathcal{S}},\hat{o}_{\mathcal{S}},\hat{u}_{\mathcal{S}}\}$ now refer to the student predictions. We derive pseudo-labels $y_{\mathcal{T}}=\{p_{\mathcal{T}},o_{\mathcal{T}},u_{\mathcal{T}}\}$ from the teacher’s predictions $\hat{y}_{\mathcal{T}}=\{\hat{p}_{\mathcal{T}},\hat{o}_{\mathcal{T}},\hat{u}_{\mathcal{T}}\}$ as follows:
$p_{\mathcal{T}}[t]=\hat{p}_{\mathcal{T}}[t]\quad;\quad o_{\mathcal{T}}[t]=\mathbbm{1}(\hat{o}_{\mathcal{T}}[t]>0);\quad u_{\mathcal{T}}[t]=\mathbbm{1}(d(p_{\mathcal{T}}[t],\hat{p}_{\mathcal{S}}[t])>\delta)$ (2)
where $t$ indexes time. The loss $\ell_{ssl}(\hat{p}_{\mathcal{S}}[t],\hat{o}_{\mathcal{S}}[t],\hat{o}_{\mathcal{S}}[t])$ for a given video frame $t$ is derived from the TAPIR loss, treating the pseudo-labels as ground-truth, and defined as:
@@
While the above formulation is well-defined, if the student and teacher both receive the same video and query point, we expect the loss to be trivially close to zero, so we apply transformations and corruptions to the student’s view. Given an input video, we create a second view by resizing each frame to a smaller resolution (varying linearly over time) and superimposing them onto a black background at a random position (also varying linearly across time); formally, this is a frame-wise axis-aligned affine transformation $\Phi$ on coordinates, applied to the pixels. We also apply $\Phi$ to the student query coordinates. We further degrade this view by applying a random JPEG degradation to make the task more difficult, before pasting it onto the black background. Both operations lose texture information; therefore, the network must learn higher-level—and possibly semantic—cues (e.g. the tip of the top left ear of the cat), rather than lower-level texture matching in order to track points correctly. We apply the inverse affine transformation $\Phi^{-1}$ to map the student’s predictions back to the original input coordinate space, before feeding these to the loss. We describe these transformations and corruptions in more detail in Appendix 0.B.1.
Second, we enforce that each track forms an equivalence class by training the model to produce the same track regardless of which point is used as a query. While we do not have access to the ground-truth trajectories to sample different query points from, we can use the teacher model’s predictions to form pairs of query points. First, we sample a query point $Q_{1}=(q_{1},t_{1})$, where $q_{1}$ is an $(x,y)$ coordinate, and $t_{1}$ is a frame index, both sampled uniformly. Then the student’s query is sampled randomly from the teacher’s trajectory, i.e. $Q_{2}=(q_{2},t_{2})\in\{(p_{\mathcal{T}}[t],t);t\textit{ s.t. }o_{\mathcal{T}}[t]=0\}$.
@@
Here, $\delta_{cycle}$ is a distance threshold hyperparameter, which we set to 4 pixels. Furthermore, we expect that the teacher’s predictions may be less accurate than the student’s for points that are closer in time to the student’s query frame than they are to the teacher’s. Therefore, we filter these points, as depicted by the grey squares/circles in Figure 2. We define a second mask as follows:
$m_{proximity}^{t}=\mathbbm{1}(|t-t_{1}|\leq|t-t_{2}|)$ (5)
Note that there is a special case when the student and teacher have the same query point: there is no longer any uncertainty regarding whether the point is on the same trajectory. These points are reliable while also being less challenging. We compromise between extremes, and have $Q_{1}=Q_{2}$ half the time (and set the masks to 1’s), and the rest of the time sample $Q_{2}$ uniformly from the visible points in the teacher prediction. The final self-supervised loss for a single trajectory is then:
$\mathcal{L}_{SSL}=\sum_{t}m_{cycle}^{t}*m_{proximity}^{t}*\ell_{ssl}^{t}$ (6)
In practice, we sample 128 query points per input video and average the loss for all of them. For a recap of the algorithm, see Appendix 0.A.
To avoid catastrophic forgetting, we continue training on the Kubric dataset with the regular supervised TAPIR loss. Our training setup follows prior work on multi-task self-supervised learning [11]: we maintain separate Adam optimizer parameters to compute separate updates for both tasks, and then apply the gradients with their own learning rates. As the self-supervised task is more expensive due to the extra forward pass, we use half the batch size for self-supervised updates, and therefore we halve the learning rate for these updates. See Appendix 0.B.2
Note that concurrent work [48] reproduces some of these decisions, including using cycle-consistency as a method of filtering the pseudo-ground-truth and using similar augmentations before passing to the student model, e.g. affine transformations. They have a filtering step based on cycle-consistency similar to $m_{cycle}$, but do not filter based on proximity like $m_{proximity}$. Rather than a student-teacher setup, they compute trajectories only once and freeze the training data, meaning that the model is permanently trained to reproduce errors in the original labeling. Furthermore, the work fine-tunes on the target dataset, meaning that new domains require a potentially large training set and a re-start of the fine-tuning; in contrast, our work demonstrates that it’s possible to train on a single large dataset that covers many domains, meaning that fine-tuning is unnecessary.

### v2

@@ Abstract
To endow models with greater understanding of physics and motion, it is useful to enable them to perceive how solid surfaces move and deform in real scenes. This can be formalized as Tracking-Any-Point (TAP), which requires the algorithm to track any point on solid surfaces in a video, potentially densely in space and time. Large-scale ground-truth training data for TAP is only available in simulation, which currently has a limited variety of objects and motion. In this work, we demonstrate how large-scale, unlabeled, uncurated real-world data can improve a TAP model with minimal architectural changes, using a self-supervised student-teacher setup. We demonstrate state-of-the-art performance on the TAP-Vid benchmark surpassing previous results by a wide margin: for example, TAP-Vid-DAVIS performance improves from 61.3% to 67.4%, and TAP-Vid-Kinetics from 57.2% to 62.5%. For visualizations, see our project webpage at https://bootstap.github.io/
@@ 1 Introduction
Despite impressive achievements in the vision and language capability of generalist AI systems, physical and spatial reasoning remain notable weaknesses of state-of-the art vision models [47, 59]. This limits their application in many domains like robotics, video generation, and 3D asset creation – all of which require an understanding of the complex motions and physical interactions in a scene. Tracking-Any-Point (TAP) [12] is a promising approach to represent precise motions in videos, and recent work has demonstrated compelling usage of TAP in robotics [62, 69, 2], 3D reconstruction [64], video generation [13], and video editing [71]. In TAP, algorithms are fed a video and a set of query points—potentially densely across the video—and must output the tracked location of these query points in the video’s other frames. If the point is not visible in a frame, the point is marked as occluded in that frame. This approach has many advantages: it is a highly general task, as correspondences for surface points are well-defined for opaque, solid surfaces, and it provides rich information about the deformation and motion of objects across long time periods.
@@
The main challenge for building TAP models, however, is the lack of training data: in the real world, we must rely on manual labeling, which is arduous and imprecise [12], or on 3D sensing [1], which is only available in limited scenarios and quantity. Thus, state-of-the-art methods have relied on synthetic data [19, 74]. In this work, however, we overcome this limitation and demonstrate that unlabeled real-world videos can be used to improve point tracking, using self-consistency as a supervisory signal. In particular, we know that when tracks are correct for a given video, then 1) spatial transformations of the video should result in an equivalent spatial transformation of the trajectories, 2) that different query points along the same trajectory should produce the same track, and 3) that non-spatial data augmentation (e.g. image compression) should not affect results. Deviations from this can be treated as an error signal for learning.
Our architecture is outlined in Figure 1. We begin with a strong “teacher” model pre-trained using supervised learning on synthetic data (in our case, a TAPIR [13] model) which serves as initialization for both a “teacher” and a “student” model. Given an unlabeled input video, we make a prediction using the teacher model, which serves as pseudo-ground-truth for the student. We then generate a second “view” of the video by applying affine transformations that vary smoothly in time, re-sampling frames to a lower resolution, and adding JPEG corruption, and padding back to the original size. We input the second view to the “student” network and use a query point sampled from the teacher’s prediction (transformed consistently with the transformation applied to the video). The student’s prediction is then transformed back into the original coordinate space. We then use a self-supervised loss (SSL) to update the student’s weights: that is, we apply TAPIR’s original loss function to the student predictions, using the teacher’s predictions as pseudo-ground-truth. The teacher’s weights are updated by using an exponential moving average (EMA) of the student’s weights. We take steps to ensure that the teacher’s predictions used for training are more likely to be accurate than the student’s: (i) the corruptions that degrade and downsample the video are only applied to the student’s inputs, (ii) we use an EMA of the student’s weights as the teacher’s weights, a common trick for stabilizing student-teacher learning [20, 58]. Co-training using this formulation on real-world videos, in addition to training on synthetic data, provides a substantial boost over prior state-of-the-art across the entire TAP-Vid benchmark.
@@
    We analyze the importance of varying model components, and show that a surprisingly simple formulation is sufficient to achieve good results.
  3. 3.
@@
  4. 4.
    We will release a model and checkpoint on GitHub, including model implementations in both JAX and PyTorch for the community to use.
@@ 2 Related Work
The ability to track densely-sampled points over long video sequences is a generic visual capability [52, 53]. Because this visual task provides a rich output that is well-defined independent of semantic or linguisitic categories (unlike classification, detection, and semantic segmentation), it is more generically useful and can support other visual capabilities like video editing [71], 3D estimation [65], object segmentation [46, 50], camera tracking [8] and even robotics [62, 69]. Point tracking has recently experienced a flurry of recent works including new datasets [12, 74, 1] and algorithms [22, 29, 13, 65, 44, 3, 43]. Current state-of-the-art works mainly train in a supervised manner, relying heavily on synthetic data [19, 74] which has a large domain gap with the real world.
@@
A final self-supervised approach is to create pseudo-ground-truth correspondences for semi-supervised training [23, 54]. Such approaches have a long history in optical flow [37, 38, 24, 45], although with mixed results, typically requiring complex training setups such as GANs [31] or connecting the student to the teacher [39] to prevent trivial solutions. They have only been applied to longer-term point tracking more recently [65, 57]. OmniMotion computes initial point tracks using RAFT [60] or TAP-Net [12] and infers a full pseudo-3D interpretation of the scene in the form of a neural network. Although this method improves point tracks compared to their initialization, it never retrains a general TAP model on the self-labeled data. Perhaps most related is Li et al. [34], which proposes a self-supervised loss based on reconstruction, in addition to supervised point tracking loss and an adversarial domain adaptation loss. The final algorithm is complex, and performs far below our work (59.8 on TAP-Vid-DAVIS $<\delta^{x}_{avg}$, versus 78.1 for our work), with the self-supervised providing a relatively small boost. Concurrent work [57], on the other hand, saves a dataset of point tracks and retrains the underlying model on them, using data augmentations similar to ours. We discuss the differences in detail in the following section, after presenting our approach.
@@ 3 Method
When developing a self-supervised training method for TAP, it is important to note that TAP has a precise, correct answer for almost every query point. This is different from typical visual self-supervised learning, where the representation can be arbitrary, as long as semantically similar images have similar representations. Supervised learning on synthetic data provides a strong initial guess in many situations, but care must be taken to ensure that the self-supervised algorithm does not find “trivial shortcuts” [11] that become self-reinforcing and harm the initialization.
@@
Figure 2: Bootstrapped training for Tracking-Any-Point. The teacher TAPIR produces a pseudo-label trajectory from query point $q_{1}$ at time $t_{1}$. Video frames undergo affine transformations $\Phi$ that vary smoothly in time and are augmented with JPEG artifacts, then fed to the student TAPIR, which predicts a trajectory from query point $\Phi(q_{2})$ at time $t_{2}$ (sampled from the teacher’s prediction, then transformed to the student video space using $\Phi$). The student trajectory is transformed back, and loss is computed against the teacher’s trajectory. To maximize the chances that we train on accurate trajectories, we remove trajectories where the student’s prediction at time $t_{1}$ is too far from the teacher query point $q_{1}$ (i.e. not cycle-consistent; light-orange disk).
@@
Our formulation relies on two facts about point tracks that are true for points on any solid, opaque surface. First, spatial transformations (e.g. affine transformations) which are applied to the video will result in equivalent spatial transformations of the point tracks (i.e. the tracks are “equivariant” under spatial transformation), while the tracks are invariant to many other factors of variation that do not move the image content (e.g. color changes, noise). Second, the algorithm should output the same track regardless of which point along the track is used as a query; mathematically, this means that each trajectory forms an equivalence class. One could imagine enforcing the desired equivariance and invariance properties using a simple Siamese-network formulation [21], where a single network is trained to output consistent predictions on two different ‘views’ of the data (i.e., augmented and transformed versions of the video and tracks). However, we find that minimizing the difference between the two outputs—and backpropping both—results in predictions degrading toward trivial solutions (e.g. over-smoothing of tracks, or tracking the image boundary instead of the image content). In fact, the model can learn to distinguish between synthetic and real data resulting in trivial solutions on the real, unlabeled data only. To prevent this, we adopt a student-teacher framework, where the student’s view of the data is made more challenging by augmentations, and the teacher does not receive gradients that may corrupt its predictions. Figure 2 shows the overall pipeline.
Loss functions.
@@
We start with a baseline TAPIR network pre-trained on Kubric following [13]. Let $\hat{y}=\{\hat{p},\hat{o},\hat{u}\}$ be the predictions: $\hat{p}\in\mathbb{R}^{T\times 2}$ is the position, $\hat{o}\in\mathbb{R}^{T}$ is an occlusion logit, and $\hat{u}\in\mathbb{R}^{T}$ is an uncertainty logit, where $T$ is the number of frames. Calling $p[t]$ and $o[t]$ the ground truth targets for frame $t$, recall that the standard TAPIR loss for a single trajectory is defined as:
@@
After pre-training, we add extra capacity to the model to absorb the extra training data: 5 layers of 2D conv-residual layers to the backbone with a channel multiplier of 4, which roughly doubles the number of parameters in the backbone (see Appendix 0.C for details). These are initialized to the identity following “zero init” [18]. Let $\hat{y}_{\mathcal{S}}=\{\hat{p}_{\mathcal{S}},\hat{o}_{\mathcal{S}},\hat{u}_{\mathcal{S}}\}$ now refer to the student predictions. We derive pseudo-labels $y_{\mathcal{T}}=\{p_{\mathcal{T}},o_{\mathcal{T}},u_{\mathcal{T}}\}$ from the teacher’s predictions $\hat{y}_{\mathcal{T}}=\{\hat{p}_{\mathcal{T}},\hat{o}_{\mathcal{T}},\hat{u}_{\mathcal{T}}\}$ as follows:
$p_{\mathcal{T}}[t]=\hat{p}_{\mathcal{T}}[t]\quad;\quad o_{\mathcal{T}}[t]=\mathbbm{1}(\hat{o}_{\mathcal{T}}[t]>0);\quad u_{\mathcal{T}}[t]=\mathbbm{1}(d(\hat{p}_{\mathcal{T}}[t],\hat{p}_{\mathcal{S}}[t])>\delta)$ (2)
where $t$ indexes time. The loss $\ell_{ssl}(\hat{p}_{\mathcal{S}}[t],\hat{o}_{\mathcal{S}}[t],\hat{u}_{\mathcal{S}}[t])$ for a given video frame $t$ is derived from the TAPIR loss, treating the pseudo-labels as ground-truth, and defined as:
@@
Video degradations.
While the above formulation is well-defined, if the student and teacher both receive the same video and query point, we expect the loss to be trivially close to zero; therefore, we apply transformations and corruptions to the student’s view of the video. Given an input video, we create a second view by resizing each frame to a smaller resolution $r$ and superimposing it onto a black background at a random position $(h,w)$ within this background. $r$ varies linearly over time, meaning that the frames gradually become larger or smaller within the fixed-size black background. Overall, the decreased resolution degrades the student view, and this increases task difficulty for the student. The location of these frames also move with time, and $(h,w)$ follows a linear trajectory within the black background. Formally, this is a frame-wise axis-aligned affine transformation $\Phi$ on coordinates, applied to the pixels. We also apply $\Phi$ to the student query coordinates. We further degrade this view by applying a random JPEG degradation to make the task more difficult, before pasting it onto the black background. Both operations lose texture information; therefore, the network must learn higher-level—and possibly semantic—cues (e.g. the tip of the top left ear of the cat), rather than lower-level texture matching in order to track points correctly. We apply the inverse affine transformation $\Phi^{-1}$ to map the student’s predictions back to the original input coordinate space, before feeding these to the loss. We describe these transformations and corruptions in more detail in Appendix 0.B.1.
Choosing the sample point.
We enforce that each trajectory forms an equivalence class by training the model to produce the same track regardless of which point is used as a query. While we do not have access to the ground-truth trajectories to sample different query points from, we can use the teacher model’s predictions to form pairs of query points. First, we sample a query point $Q_{1}=(q_{1},t_{1})$, where $q_{1}$ is an $(x,y)$ coordinate, and $t_{1}$ is a frame index, both sampled uniformly. Then the student’s query is sampled randomly from the teacher’s trajectory, i.e. $Q_{2}=(q_{2},t_{2})\in\{(p_{\mathcal{T}}[t],t);t\textit{ s.t. }o_{\mathcal{T}}[t]=0\}$.
@@
Here, $\delta_{cycle}$ is a distance threshold hyperparameter, which we set to 4 pixels.
Note that there is a special case when the student and teacher have the same query point: there is no longer any uncertainty regarding whether the point is on the same trajectory. These points are reliable while also being less challenging. We compromise between extremes, and sample $Q_{1}=Q_{2}$ with probability 0.5, and sample with equal probabilities from the remaining visible points in the teacher prediction. The final self-supervised loss for a single trajectory is then:
$\mathcal{L}_{SSL}=\sum_{t}m_{cycle}^{t}*\ell_{ssl}^{t}$ (5)
In practice, we sample 128 query points per input video and average the loss for all of them. We provide pseudocode for the algorithm in Appendix 0.A.
To avoid catastrophic forgetting, we continue training on the Kubric dataset with the regular supervised TAPIR loss. Our training setup follows prior work on multi-task self-supervised learning [14]: we maintain separate Adam optimizer parameters to compute separate updates for both tasks, and then apply the gradients with their own learning rates. As the self-supervised task is more expensive due to the extra forward pass, we use half the batch size for self-supervised updates, and therefore we halve the learning rate for these updates. See Appendix 0.B.2 for more details.
Differences between our approach and [57]. Concurrent work reproduces some of these decisions, including using cycle-consistency as a method of filtering and using affine transformations when augmenting the student view. However, there are a few key differences. First, rather than a student-teacher setup, they compute trajectories only once and freeze the training data, meaning that the model is permanently trained to reproduce errors in the original labeling. Furthermore, the work fine-tunes on the target dataset, meaning that transfer to a new domain may require a large training set in that domain on which to fine-tune; in contrast, our work demonstrates that it’s possible to train on a single large dataset that covers many domains, meaning that fine-tuning is unnecessary.